1. A firm manufactures two products; the net profit on product 1 is Rupees 3 per unit and Rupees 5 per unit on product 2. The manufacturing process is such that each product has to be processed in two departments D1 and D2. Each unit of product1 requires processing for 1 minute at D1 and 3 minutes at D2; each unit of product 2 requires processing for 2 minutes at D1 and 2 minutes at D2. Machine time available per day is 860 minutes at D1 and 1200 minutes at D2. How much of product 1 and 2 should be produced every day so that total profit is maximum. (solve with graphical method)

1
Expert's answer
2021-11-02T13:01:33-0400

"3\cdot x_1+5\cdot x_2\rarr max"

"x_1+2\cdot x_2\le800;\\
3\cdot x_1+2\cdot x_2\le1200;\\"

"x_1\ge0,x_2\ge0"


Extreme Point Coordinates:

"(400,0)" : "z=1200"

"(200,300)" : "z=2100"

"(0,400)" : "z=2000"


optimal solution:

 x1=200,x2=300 and max Z=2100.